I am doing the substitution of equations method. y=5-4x,2x-3y=13. When I rewrite as 5-4x+2x-3y=13.Would I first have to subtract 5 from 4x before mulitiplying the x's.
Answer by stanbon(75887)

y=5-4x
2x-3y=13
-----------
Substitute for "y" and solve for "x":
2x -3(5-4x) = 13
2x - 15 + 12x = 13
14x = 28
x = 2
-------
Solve for "y" using y = 5-4x
y = 5 - 4*2
y = -3
============
Cheers,
Stan H.
============